OpenStudy (anonymous):

which is true statement about functions? A. if f and g are functnions then (g o f)(x)=(f o g)(x) B. if a and b are matrices, then ab can always be computed. C. if f and g are functions then (f+g)(-2)=(g+f)(-2) D. if f is a function then f(2+h)=f(2)+f(h)

jimthompson5910 (jim_thompson5910):

which one do you think it is?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

nvm d

OpenStudy (anonymous):

lol flutter, it was c... :C

OpenStudy (anonymous):

now, why didn't d work though o.o

jimthompson5910 (jim_thompson5910):

Example of why D doesn't work f(x) = x^2 f(2+h) = (2+h)^2 f(2+h) = 4 + 2h + h^2 ------------------------ f(2) = 2^2 f(2) = 4 ------------------------ f(h) = h^2 =========================== So f(2+h) = f(2)+f(h) becomes 4 + 2h + h^2 = 4 + h^2 which is only true when h = 0

OpenStudy (anonymous):

hmm, I forgot about the 2h... I'm not sure where the 2h came from though. did you distribute ^2 and then aftewards add 2+h

jimthompson5910 (jim_thompson5910):

(2+h)^2 = (2+h)(2+h) (2+h)^2 = 2(2+h) + h(2+h) (2+h)^2 = 2*2+2h + 2h+h*h (2+h)^2 = 4+2h + 2h+h^2 (2+h)^2 = 4+4h+h^2 hmm sry about that, made a typo and it should be (2+h)^2 = 4+4h+h^2 and not 4+2h+h^2, my bad but it still holds true why D doesn't work